Solve the equation algebraically: log (base 1.2) x = 1-log (base 1.2) (x-0.2)
Answer by Alan3354(69443)

logs base 1.2
----
log(x) = 1-log(x-0.2)
log(x) = log(1.2) - log(x-0.2)
log(x) + log(x-0.2) = log(1.2)
log(x*(x-0.2)) = log(1.2)
Since logs have the same base:
---
x*(x-0.2) = 1.2
Solve the quadratic.
Check the solutions to avoid logs of negatives.